# Correlations for aerodynamic force coefficients of non-spherical particles in compressible flows

### v1

@@ 4.2.2 Oblate spheroid
The results of the evolution of the lift coefficient as a function of the angle of attack for the oblate spheroid, see the centre of Figure 15, show that the values for the lift coefficients are almost symmetric with respect to the angle of attack $\alpha=45\degree$. This trend indicates that the shift in the angle of attack at which the maximum lift coefficient occurs is less influenced by the increase in the particle Reynolds number than for the rod-like and the prolate spheroid particles. The specific evolution of the lift coefficient that varies largely from the other shapes confirms the necessity to have a unique correlation to predict the lift coefficient, which relies on the modification of the exponent in the cosine-sine multiplication, realized by the exponents $\vartheta_{4}$ and $\vartheta_{5}$ in the cosine-sine multiplication. In general, the correlation agrees well with the PR-DNS results, as the maximum and median prediction errors between the correlation and the PR-DNS are 9.91% and 1.02%, respectively.
@@ 4.2.3 Rod-like particle
$\begin{split}\mathcal{F}_{\mathrm{L}}^{\mathrm{Ma}}\left(\mathrm{Re_{p}},\alpha,\mathrm{Ma}\right)=&\frac{\vartheta_{1}\log\left(\mathrm{Re_{p}}\right)^{\vartheta_{2}}}{\left[1+\exp\left(\vartheta_{3}\cdot(\mathrm{Ma}-\vartheta_{4})\right)\right]}\cos\left(\Psi_{1}\left(\alpha,\mathrm{Re_{p}}\right)\right)\sin\left(\Psi_{1}\left(\alpha,\mathrm{Re_{p}}\right)\right)\\ &+\vartheta_{5}\exp\left(\frac{-(\mathrm{Ma}-\vartheta_{6})^{2}}{\vartheta_{7}}\right)\left[\cos\left(\Psi_{2}\left(\alpha,\mathrm{Re_{p}}\right)\right)\sin\left(\Psi_{2}\left(\alpha,\mathrm{Re_{p}}\right)\right)\right]^{\vartheta_{8}}\,.\end{split}$ (18)
@@
$\Psi_{2}=\cfrac{\pi}{2}\left(\alpha\cfrac{2}{\pi}\right)^{1+48.896\log{\left(\mathrm{Re_{p}}\right)}^{-3.182}}$ (20)
@@
The PR-DNS results along with the correlation for the prediction of the rod-like particle lift coefficient are shown in the right plot of Figure 15, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The correlation accurately recovers the evolution of the lift coefficient in all regimes, for all particle Reynolds numbers and angles of attack. The maximum and median prediction errors between the correlation and the PR-DNS are 10.88% and 0.62%, respectively.
@@ 4.3 Torque coefficient
$\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}=\left[\frac{f_{1}\left(\mathrm{Re_{p}},\vartheta\right)}{1+\exp\left(\mathrm{Ma}-f_{2}\left(\mathrm{Re_{p}},\vartheta\right)\right)}+f_{3}\left(\mathrm{Re_{p}},\vartheta\right)\exp\left(f_{4}\left(\mathrm{Re_{p}},\vartheta\right)\right)\right]\cos\left(\Psi\left(\alpha,\mathrm{Re_{p}}\right)\right)\sin\left(\Psi\left(\alpha,\mathrm{Re_{p}}\right)\right)\,,$ (22)
where the sigmoid and the Gaussian terms are multiplied by a shifted cosine-sine expression, because the torque coefficient is no longer largest for a particle with angle of attack $\alpha=45\degree$, but for slightly larger angles of attack [15, 38, 9]. The expressions of the shape-dependent functions $\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}$ for each particle are given in the dedicated sections, along with the associated prediction errors. The fitting coefficients, $\vartheta_{i}$, for the torque coefficients of all three considered particle shapes are given in Table 3.
@@
$\vartheta_{1}$ $\vartheta_{2}$ $\vartheta_{3}$ $\vartheta_{4}$ $\vartheta_{5}$ $\vartheta_{6}$ $\vartheta_{7}$ $\vartheta_{8}$ $\vartheta_{9}$
Eq. (23) 0.219 0.200 12.5 0.560 0.120 1.09 0.150 600 9.48 $\times$ $10^{-2}$
Eq. (25) 0.810 28.8 0.823 0.442 3.87 $\times$ $10^{-2}$ 1.31 0.450 4093 -
Eq. (27) 0.157 0.200 10.1 1.46 0.355 1.17 0.150 600 -
4.3.1 Prolate spheroid
To consider compressible effects in the prediction of the torque coefficient of an isolated prolate spheroid particle, the shape-dependent function, $\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}$, in Eq. (21), is given by
$\begin{split}\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}=&\left[\frac{\vartheta_{1}\mathrm{Re_{p}}^{\vartheta_{2}}}{\left[1+\exp\left(-\vartheta_{3}(\mathrm{Ma}-\vartheta_{4}\mathrm{Re_{p}}^{\vartheta_{9}})\right)\right]}+\vartheta_{5}\mathrm{Re_{p}}^{\vartheta_{2}}\exp\left(-\frac{(\mathrm{Ma}-\vartheta_{6})^{2}}{\vartheta_{7}}-\frac{\mathrm{Re_{p}}}{\vartheta_{8}}\right)\right]\\ &\times\cos\left(\Psi(\alpha,\mathrm{Re_{p}})\right)\sin\left(\Psi(\alpha,\mathrm{Re_{p}})\right)\,,\end{split}$ (23)
with the function to shift the maximum value of the cosine-sine product given as
$\Psi=\cfrac{\pi}{2}\left(\alpha\cfrac{2}{\pi}\right)^{1.438+0.114\log{\left(\mathrm{Re_{p}}^{1/6}\right)}}$ (24)
The results of the PR-DNS along with the correlation for the torque coefficient of the prolate spheroid are shown in the left of Figure 16, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The change from the subsonic to the supersonic flow regimes is accurately recovered for all considered angles of attack and particle Reynolds numbers by the sigmoid term in Eq. (23). The shift in the angle of attack at which the maximum torque coefficient occurs is also accurately recovered by the cosine-sine product. This can be observed in the left of Figure 16, where the torque coefficient at $\alpha=75\degree$ is lower than the torque coefficient at $\alpha=15\degree$. Also, the smooth increase of the torque coefficient as a function of the Mach number in the transonic regime is recovered accurately for all considered cases. This leads to maximum and median prediction errors between the correlation and the PR-DNS of 5.42% and 0.22%, respectively.
4.3.2 Oblate spheroid
For the oblate spheroid, the shape-dependent function, $\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}$, in Eq. (21), is given by
$\begin{split}\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}=&\left[\frac{\vartheta_{1}}{\left[1+\exp\left(-\vartheta_{2}(\mathrm{Ma}-\vartheta_{3})\right)\right]}+\vartheta_{4}\mathrm{Re_{p}}^{\vartheta_{5}}\exp\left(-\frac{(\mathrm{Ma}-\vartheta_{6})^{2}}{\vartheta_{7}}-\frac{\mathrm{Re_{p}}}{\vartheta_{8}}\right)\right]\\ &\times\cos(\Psi(\alpha))\sin(\Psi(\alpha))\,,\end{split}$ (25)
with the function to shift the maximum value of the cosine-sine product given as
$\Psi=\cfrac{\pi}{2}\left(\alpha\cfrac{2}{\pi}\right)^{1.531}$ (26)
As for the prolate spheroid, the model prediction for the torque coefficient is shown with the PR-DNS results in the centre of Figure 16. For the oblate spheroid, the evolution of the torque coefficient in the transonic regime does not systematically follow the bell-shape form observed for the prolate spheroid. For instance, for a particle with angle of attack of $\alpha=30\degree$, at $\mathrm{Re_{p}}=100$, the evolution of the torque coefficient deviates from the bell-shape evolution observed for other angles of attack. This is also observed at $\mathrm{Re_{p}}=300$ for a particle with angle of attack $\alpha=60\degree$. Since this behaviour is not consistent over all angles of attack, a Gaussian function term is retained in Eq. (25). This Gaussian function term increases the maximum prediction error between the model fit and the PR-DNS to 9.11%, which is significantly higher compared to the prolate spheroid results, but allows to preserve a relatively good agreement between the model fit and the PR-DNS since the median prediction error remains low at 0.90%.
4.3.3 Rod-like particle
The shape-dependent function, $\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}$, in Eq. (21), is given for the rod-like particle by
$\begin{split}\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}=&\left[\frac{\vartheta_{1}\mathrm{Re_{p}}^{\vartheta_{2}}}{\left[1+\exp\left(-\vartheta_{3}(\mathrm{Ma}-\vartheta_{4})\right)\right]}+\vartheta_{5}\mathrm{Re_{p}}^{\vartheta_{2}}\exp\left(-\frac{(\mathrm{Ma}-\vartheta_{6})^{2}}{\vartheta_{7}}-\frac{\mathrm{Re_{p}}}{\vartheta_{8}}\right)\right]\\ &\times\cos\left(\Psi(\alpha,\mathrm{Re_{p}})\right)\sin\left(\Psi(\alpha,\mathrm{Re_{p}})\right)\,,\end{split}$ (27)
where the expression to shift the maximum value of the cosine-sine product is very similar to the expression for the prolate spheroid, given by
$\Psi=\cfrac{\pi}{2}\left(\alpha\cfrac{2}{\pi}\right)^{1.676+0.176\log{\left(\mathrm{Re_{p}}^{1/6}\right)}}\,$ (28)
The model prediction for the torque coefficient is shown against the PR-DNS results in the right of Figure 16. As for the torque coefficient of the prolate spheroid, that has a similar shape, an excellent agreement is observed for all considered cases, from the subsonic to the supersonic flow regime, with relatively low maximum and median prediction errors between the correlation and the PR-DNS of 5.44% and 0.22%, respectively.
@@ 5 Conclusions
In this study, we conduct particle-resolved direct numerical simulations (PR-DNS) with three-dimensional body-fitted hexahedral grids to investigate the aerodynamic force and torque coefficients of particles with different shapes, using a finite-volume-based solver for compressible flows. The employed computational domain is spherical, centred around the particle, facilitating the rotation of the particle without changing the computational grid. This computational domain is represented by non-uniform and body-conforming grid, refined near the particle, and with adaptive mesh refinement (AMR) based on the local pressure gradient, the vorticity, and the non-dimensional wall distance. The PR-DNS cover three particle shapes, namely a prolate spheroid with an aspect ratio of 5/2, an oblate spheroid with an aspect ratio of 5, and a rod-like particle with an aspect ratio of 5. The simulations are carried out at a reference temperature of 300 K and a reference pressure of $1.0\times 10^{5}$ Pa, maintaining a Prandtl number of 0.71.
@@
This study has examined the behaviour of non-spherical particles in compressible flows, focusing on the steady-state velocity as well as the aerodynamic force coefficients (drag, lift, and pitching torque) for the three considered particle shapes at Mach numbers ranging from 0.3 to 2.0, angles of attack from 0°to 90°, and particle Reynolds numbers between 100 and 300. The bow shock developing in the supersonic flow regime in front of the particle is largest for the oblate spheroid and smallest for the prolate spheroid. In the transonic regime, the oblate spheroid exhibits unstable wake behaviour when the angle of attack is larger than $0\degree$. The drag coefficients increase with the angle of attack and are highest for the oblate spheroid. All shapes show a rapid increase in drag in the transonic regime, followed by a decrease in the supersonic regime. The lift and pitching torque coefficients follow similar trends, with the oblate spheroid having the highest values. We can, therefore conclude that the shape of the particles impacts flow behaviour and aerodynamic forces significantly, with oblate spheroids showing the most pronounced effects due to their geometric characteristics.
@@
The correlations proposed in this work are valid for particle Reynolds numbers between 100 and 300, Mach numbers between 0 and 2, and angles of attack ranging from $0\degree$ to $90\degree$. The correlations accurately recover the change in the evolution of the aerodynamic force and torque coefficients in both transonic and supersonic regimes, for all considered particle Reynolds numbers and angles of attack. The largest maximum error of the correlations for the drag coefficient is 8.65% for the prolate spheroid, while the median error for the three considered particle shapes is approximately 1%. The proposed correlations for the lift coefficient have a slightly larger maximum error, with the rod-like particle correlation having the highest maximum error of 10.88%. However, the median error of the correlations for the lift coefficient is also approximately 1%. For the torque coefficient, the correlation of the oblate spheroid has the largest maximum error with 9.11%, whereas the maximum errors of the correlations for the other two considered shapes are significantly lower, with a maximum error of around 5%. Again, the median error of the torque correlations for all three particles shapes is approximately 1%.
@@
The data that support the findings of this study are reproducible and files to regenerate the data as well as an executable to implement the correlations to predict the drag, lift, and torque coefficients are openly available in the repository with DOI 10.5281/zenodo.13364263

### v2

@@ 4.2.2 Oblate spheroid
The results of the evolution of the lift coefficient as a function of the angle of attack for the oblate spheroid, see the centre of Figure 12, show that the values for the lift coefficients are almost symmetric with respect to the angle of attack $\alpha=45\degree$. This trend indicates that the shift in the angle of attack at which the maximum lift coefficient occurs is less influenced by the increase in the particle Reynolds number than for the rod-like and the prolate spheroid particles. The specific evolution of the lift coefficient that varies from the other shapes confirms the necessity to have a unique correlation to predict the lift coefficient. This unique modification relies on the change of the exponent in the cosine-sine multiplication, realized by the exponents $\vartheta_{4}$ and $\vartheta_{5}$. In general, the correlation agrees well with the PR-DNS results, as the maximum and median prediction errors between the correlation and the PR-DNS are 13.66% and 0.19%, respectively.
@@ 4.2.3 Rod-like particle
$\begin{split}\mathcal{F}_{\mathrm{L}}^{\mathrm{Ma}}\left(\mathrm{Re_{p}},\alpha,\mathrm{Ma}\right)=&\frac{{{\left(\vartheta_{1}/\log\left(\mathrm{Re_{p}}\right)\right)}}^{\vartheta_{2}}}{\left[1+\exp\left(\vartheta_{3}\cdot(\mathrm{Ma}-\vartheta_{4})\right)\right]}\cos\left(\Psi_{1}\left(\alpha,\mathrm{Re_{p}}\right)\right)\sin\left(\Psi_{1}\left(\alpha,\mathrm{Re_{p}}\right)\right)\\ &+\vartheta_{5}\exp\left(\frac{-(\mathrm{Ma}-\vartheta_{6})^{2}}{\vartheta_{7}}\right)\left[\cos\left(\Psi_{2}\left(\alpha,\mathrm{Re_{p}}\right)\right)\sin\left(\Psi_{2}\left(\alpha,\mathrm{Re_{p}}\right)\right)\right]^{\vartheta_{8}}\,.\end{split}$ (19)
@@
$\Psi_{2}=\cfrac{\pi}{2}\left(\alpha\cfrac{2}{\pi}\right)^{1+{{1630.79\left(1/\log{\left(\mathrm{Re_{p}}\right)}\right)^{6.00}}}}$ (21)
@@
The PR-DNS results along with the correlation for the prediction of the rod-like particle lift coefficient are shown in the right plot of Figure 12, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The correlation accurately recovers the evolution of the lift coefficient in all regimes, for all particle Reynolds numbers and angles of attack. The maximum and median prediction errors between the correlation and the PR-DNS are 7.97% and 0.15%, respectively.
@@ 4.3 Torque coefficient
$\begin{split}\mathcal{F}_{\mathrm{M}}^{\mathrm{Ma}}&=\frac{\vartheta_{1}\mathrm{Re_{p}}+\vartheta_{2}\left(\mathrm{Re_{p}}/\vartheta_{3}\right)^{\vartheta_{4}}}{1+\exp\left(-\vartheta_{5}\left(\mathrm{Ma}-\vartheta_{6}\right)\right)}\cdot\left(\cos\left(\Psi_{1}\right)\sin\left(\Psi_{1}\right)\right)^{\vartheta_{7}}\\ &+\left(\frac{\vartheta_{8}\alpha}{\vartheta_{10}}\right)^{\vartheta_{9}}\cdot\left(\frac{\vartheta_{11}\mathrm{Re_{p}}}{\vartheta_{13}}\right)^{\vartheta_{12}}\cdot\exp\left(\frac{-\left(\mathrm{Ma}-\vartheta_{14}\right)^{2}}{\vartheta_{15}}\right)\cdot\left(\cos\left(\Psi_{2}\right)\sin\left(\Psi_{2}\right)\right)^{\vartheta_{16}},\end{split}$ (23)
with the functions to shift the maximum value of the cosine-sine product given as
${{\Psi_{1}=\cfrac{\pi}{2}\left(\alpha\cfrac{2}{\pi}\right)^{\vartheta_{17}},}}$ (24)
and
${{\Psi_{2}=\cfrac{\pi}{2}\left(\alpha\cfrac{2}{\pi}\right)^{\vartheta_{18}}.}}$ (25)
As for the drag and lift correlations, the correlation to predict the torque coefficient is split in two main terms: the sigmoid term to model the change in the torque coefficient from subsonic to supersonice regimes, and a Gaussian function to model the smooth evolution of the coefficient from the subsonic to the supersonic flow regime. Both terms are multiplied by a shifted cosine-sine factor so that the torque coefficients remain zero for the angles of attack $\alpha=0\degree$ and $\alpha=90\degree$, and to model the shift in the maximum torque coefficient. The fitting coefficients, $\vartheta_{i}$, for the torque coefficients of all three considered particle shapes are given in Table 3. The results of the PR-DNS along with the correlation for the torque coefficient of the prolate spheroid are shown in the left of Figure 13, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The change from the subsonic to the supersonic flow regimes is accurately recovered for all considered angles of attack and particle Reynolds numbers by the sigmoid term, leading to maximum and median prediction errors between the correlation and the PR-DNS of 6.88% and 0.44%, respectively. As for the prolate spheroid, the model prediction for the torque coefficient of the oblate spheroid is shown with the PR-DNS results in the centre of Figure 13. The evolution of the torque coefficient of the oblate spheroid is very similar compared with the prolate spheroid and the correlation accurately recovers the smooth transition from the subsonic to the supersonic flow regime, leading to maximum and median prediction errors between the correlation and the PR-DNS of 12.96% and 1.40%, respectively. For the rod-like particle, the model prediction for the torque coefficient is shown against the PR-DNS results in the right of Figure 13. In this case, the transition from subsonic to supersonic flow regime is much more significant compared to the other two particle shapes and the PR-DNS results in the transition region are more narrow, which makes it more difficult to find an accurate model prediction. Hence, the maximum prediction error between the correlation and the PR-DNS is 22.40% because of a single outlier, but the median prediction error is 0.18%, which shows that the correlation is over all in an excellent agreement with the PR-DNS results.
@@
Prolate Oblate Rod-like
$\vartheta_{1}$ 2.28 ×$\times 10^{-4}$4.43 ×$\times 10^{-3}$2.32 ×$\times 10^{-5}$
$\vartheta_{2}$0.444 -0.759 -1.15
$\vartheta_{3}$ 157 203 127
$\vartheta_{4}$0.223 -1.03 -0.134
$\vartheta_{5}$ 15.1 5.93 12.3
$\vartheta_{6}$ 1.01 1.22 1.03
$\vartheta_{7}$ 0.993 1.11 1.04
$\vartheta_{8}$ 4.96 ×$\times 10^{-2}$ 3.18 ×$\times 10^{-9}$ 13.8
$\vartheta_{9}$ 1.44 8.95 ×$\times 10^{-2}$ 1.88
$\vartheta_{10}$ 7.42 ×$\times 10^{-3}$ 3.59 ×$\times 10^{-2}$ 3.64
$\vartheta_{11}$ 3.81 ×$\times 10^{-2}$ 3.89 0.520
$\vartheta_{12}$ 1.68 ×$\times 10^{-2}$ 0.569 0.158
$\vartheta_{13}$ 5.86 ×$\times 10^{-9}$ 10.8 1.22 ×$\times 10^{-6}$
$\vartheta_{14}$ 1.10 1.98 1.18
$\vartheta_{15}$ 9.29 ×$\times 10^{-3}$ 1.04 3.71 ×$\times 10^{-3}$
$\vartheta_{16}$ 1.05 2.50 1.16
$\vartheta_{17}$ 0.164 1.17 0.157
$\vartheta_{18}$ 1.04 0.809 1.32
@@ 5 Conclusions
In this study, we have presented particle-resolved direct numerical simulations (PR-DNS) with three-dimensional body-fitted hexahedral meshes to investigate the aerodynamic force and torque coefficients of particles with different shapes, using a finite-volume-based solver for compressible flows. The employed computational domain is spherical, centred around the particle, facilitating the rotation of the particle without changing the computational mesh. This computational domain is represented by non-uniform and body-conforming meshes, refined near the particle. PR-DNS are carried out for three particle shapes, namely a prolate spheroid with an aspect ratio of 5/2, an oblate spheroid with an aspect ratio of 5, and a rod-like particle with an aspect ratio of 5. The simulations are carried out at a reference temperature of 300 K and a reference pressure of $10^{5}$ Pa, maintaining a Prandtl number of 0.71.
@@
This study has examined the behaviour of non-spherical particles in compressible flows, focusing on the steady-state velocity as well as the aerodynamic force coefficients (drag, lift, and torque) for the three considered particle shapes at Mach numbers ranging from 0.3 to 2.0, angles of attack from 0°to 90°, and particle Reynolds numbers between 100 and 300. The bow shock developing in the supersonic flow regime in front of the particle is largest for the oblate spheroid and smallest for the prolate spheroid. In the transonic regime, the oblate spheroid exhibits an unstable wake behaviour when the angle of attack is larger than $0\degree$. The drag coefficients increase with the angle of attack and are highest for the oblate spheroid. All shapes show a rapid increase in drag in the transonic regime, followed by a decrease in the supersonic regime. The lift and torque coefficients follow similar trends, with the oblate spheroid having the highest values. We can, therefore, conclude that the shape of the particles impacts the flow behaviour and aerodynamic forces significantly, with oblate spheroids showing the most pronounced effects due to their geometric characteristics.
@@
The correlations proposed in this work are valid for particle Reynolds numbers between 100 and 300, Mach numbers between 0 and 2, and angles of attack ranging from $0\degree$ to $90\degree$. The correlations accurately recover the change in the evolution of the aerodynamic force and torque coefficients in both transonic and supersonic regimes, for all considered particle Reynolds numbers and angles of attack. The largest maximum error of the correlations for the drag coefficient is 10.19% for the prolate spheroid, while the median error for the three considered particle shapes is approximately 1-2%. The proposed correlations for the lift coefficient have a slightly larger maximum error, with the prolate particle correlation having the highest maximum error of 13.93%. However, the median error of the correlations for the lift coefficient is below 1% which highlights the overall excellent accuracy. For the torque coefficient, the correlation of the rod-like particle has the largest maximum error with 22.4%. Again, the median error of the torque correlations for all three particles shapes is approximately 1%, which indicates that the relatively high maximum error comes from a single outlier, but the overall correlations are in excellent agreement.
@@
While our current study focuses on oblate spheroids, prolate spheroids, and rod-like particles, these shapes serve as idealized cases that span a range of aspect ratios and flow characteristics. For particles with shapes that deviate from these categories, various approaches could be considered. One possibility could be the interpolation between the correlations. If the particle shape is approximately spheroidal or rod-like but does not fit perfectly within our definitions, users might interpolate between the aerodynamic force coefficients for oblate, prolate, and rod-like shapes based on the particle aspect ratio and geometry. For irregular or non-standard shapes, the methodology used in this study (e.g., using numerical simulations or experimental data) could be extended to derive additional correlations for those shapes. Alternatively, a shape factor or a drag coefficient correction based on empirical or computational approaches could be applied. Users may also combine our correlations with existing generalized drag or force models that account for irregular particle shapes through empirical shape factors or coefficients.
@@
The data that support the findings of this study are reproducible and files to regenerate the data as well as an executable to implement the correlations to predict the drag, lift, and torque coefficients are openly available in the repository with DOI 10.5281/zenodo.14283844